OpenStudy (lisa123):

Factor (2x+1)^2-4(2x+1)

OpenStudy (freckles):

do you know how to factor p^2-5p?

OpenStudy (lisa123):

yes @freckles p(p-5)

OpenStudy (badhi):

then think that \(p = (2x+1)\) and look at the question again :D

OpenStudy (phebe):

>>>>>>>>>>>>>>>>>>>>>>(2 x+1) (2 x-3) <<<<<<<<<<<<<<<<<<<

OpenStudy (phebe):

WELCOME TO OPENSTUDY!!

OpenStudy (lisa123):

how did you get that @Phebe

OpenStudy (phebe):

Factor

OpenStudy (disqin):

\[(2x+1)^2-4(2x+1)=(2x+1)\cdot [(2x+1)-4]\] \[=(2x+1)\cdot(2x-3)\]

OpenStudy (lisa123):

what happened to the ^2

OpenStudy (phebe):

Yep thats the right answer @disqin

OpenStudy (disqin):

@Phebe :-)

OpenStudy (lisa123):

Oh I understand now. Thanks!